An increase in investment demanded will shift the aggregate demand curve to the right by an amount equal to the

A money multiplier.

B investment spending multiplier.

C change in investment.

D marginal propensity to consume.
Answer #1

Answer B. Investment spending multiplier

Reason- Investment spending multiplier= ∆Y/∆I= 1/(1-MPC)=1/MPS

If investment demanded increases by $2 billion, multipier is 3, aggregate demand will rise by 3*$2 billion= $6 billion